OpenStudy (anonymous):

What is the product in lowest terms? -6/11 x 3/4 A. -9/22 B. -6/13 C. 9/22 D. 6/13

OpenStudy (anonymous):

It's the same thing principle: \[\frac{ -6 }{ 11 }*\frac{ 3 }{ 4 }=\frac{ -6*3 }{ 11*4 }=\frac{ ? }{ ? }\]

OpenStudy (anonymous):

-18 and 44

OpenStudy (anonymous):

you might want to check the solutions that your wrote though, or the question, cause the correct answer to the question you wrote isn't there

OpenStudy (anonymous):

but yes that is right. So what divides both -18 and 44?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

okay, whatever divides 18, also divides -18 so what divides 18 and 44 (hint they are both even numbers)

OpenStudy (anonymous):

good you've edited the answers :)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

not quite, 6 doesn't divide into 44. (it divides into 42 and 48 but not 44)

OpenStudy (anonymous):

11

OpenStudy (anonymous):

nope, because 11 doesn't divide into 18. (hint it's quite a small number this time)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yes! so if you have: \[\frac{ -18\div2 }{ 44\div2 }=\frac{ ? }{ ? }\] what do you get?

OpenStudy (anonymous):

9 and 22

OpenStudy (anonymous):

exactly. though the 9 will be negative (as it was -18)